OpenStudy (logan13):

Subtract. –7 – 12 A. –19 B. –5 C. 5 D. 19

OpenStudy (logan13):

these are negatives

OpenStudy (ismugglehuggles):

D

OpenStudy (anonymous):

I think A cause when negative is first you changed the subtract to add

OpenStudy (anonymous):

@logan13

OpenStudy (ismugglehuggles):

if you subtract a negative from a positive , you get a positive .

OpenStudy (anonymous):

no check on your calculator

OpenStudy (anonymous):

@logan13 D is wrong

OpenStudy (anonymous):

it A

OpenStudy (anonymous):

@Firejay5

OpenStudy (firejay5):

the answer is A. -19, because -7 - 12 = -19, *use a calculator on the computer if needed*

OpenStudy (firejay5):

Rule: 1. If you subtract a negative from a negative you add them since its a double negative 2. If you subtract a negative from a positive you'll get a negative 3. If you add a negative to a positive then your answer depends on what the prolem is.

OpenStudy (ismugglehuggles):

If you subtract a negative number from a positive number the answer would be positive , because you are taking the negative away . (-7)-(+12) = x (+7)+(+12) = 19

OpenStudy (firejay5):

incorrect @iSmuggleHuggles type -7 - 12 into your calculator onto the computer, because haha chances are the answer is -19 and not 19, don't test me and what you said is wrong if you subtract a negative from a positive the answer is negative, due to the fact that they are both different signs = -7 - (+12): Read this for more information: http://www.mathsisfun.com/positive-negative-integers.html

OpenStudy (anonymous):

Nice try Smuggle

OpenStudy (anonymous):

@iSmuggleHuggles